# Enhancing Text Editing for Grammatical Error Correction: Arabic as a Case Study

### v1

@@ Appendix A Hyperparameters
We use Hugging Face’s Transformers to build our edit taggers. Models trained on QALB-2014 or MADAR CODA are fine-tuned for 10 epochs using a learning rate of 5e-5, a batch size of 32, a maximum sequence length of 512, and a seed of 42 on a single A100 GPU. For models trained on QALB-2014 with the tenfold upsampled ZAEBUC, we use the same hyperparameters but run training for 15 epochs. At the end of the fine-tuning, we pick the best checkpoint based on the performance on the Dev sets by using the M2 scorer.
Appendix B Edits Coverage
@@
Table 8: Edit statistics on QALB-2014, ZAEBUC and MADAR CODA. Input is the input unit representation (word or subword). Comp. indicates whether the edit is compressed. Subset specifies whether the edits capture all errors, punctuation-only errors (Pnx), or non-punctuation errors (NoPnx). Edits represents the total number of unique edits in the training set of each dataset. OOVis the percentage of out-of-vocabulary edits (non-unique) in the Dev set of each dataset.
@@ Appendix C Edit Tagging Results
Table 9: MSA and DA GEC results on the Dev sets of QALB-2014, ZAEBUC, and MADAR CODA in terms of Precision (P), Recall (R), F1, and F0.5. Input is the input unit representation (word or subword). Comp. indicates whether the edit is compressed. Subset specifies whether the edits capture all errors, punctuation-only errors (Pnx), or non-punctuation errors (NoPnx). NoPnx models are evaluated after removing punctuation, while Pnx models are evaluated on a version of the Dev set where all non-punctuation errors are corrected. Pruning experiments were conducted using the top two models (AraBERTv02 and CAMeLBERT), while punctuation segregation experiments used the best model (AraBERTv02). Best All results are in bold. Best NoPnx and Pnx results are underlined.
@@ Appendix D LLMs Results
Table 10: LLMs results on MSA and DA GEC on the Dev sets of QALB-2014, ZAEBUC, and MADAR CODA in terms of Precision (P), Recall (R), F1, and F0.5. P-Lang is the prompt language either in English (EN) or Arabic (AR). Best average F0.5 results for each LLM are underlined. Best overall results are in bold.
Appendix E Prompts

### v2

@@ 6 Conclusion and Future Work
Acknowledgments
We thank Ted Briscoe for helpful discussions and constructive feedback. We also acknowledge the support of the High Performance Computing Center at New York University Abu Dhabi.
@@ Appendix A Hyperparameters
We use Hugging Face’s Transformers to build our edit taggers. Models trained on QALB-2014 or MADAR CODA are fine-tuned for 10 epochs using a learning rate of 5e-5, a batch size of 32, a maximum sequence length of 512, and a seed of 42 on a single A100 GPU. For models trained on QALB-2014 with the tenfold upsampled ZAEBUC, we use the same hyperparameters but run training for 15 epochs. At the end of fine-tuning, we pick the best checkpoint based on the performance on the Dev sets by using the M2 scorer.
Appendix B Edit Coverage
@@
Table 9: Edit statistics on QALB-2014, ZAEBUC and MADAR CODA. Input is the input unit (word or subword). Comp. indicates whether the edit is compressed. Subset specifies whether the edits capture all errors, punctuation-only errors (Pnx), or non-punctuation errors (NoPnx). Edits represents the total number of unique edits in the training set of each dataset. OOVis the percentage of out-of-vocabulary edits (non-unique) in the Dev set of each dataset.
@@ Appendix C Edit Tagging Results
Table 10: MSA and DA GEC results on the Dev sets of QALB-2014, ZAEBUC, and MADAR CODA. Input is the input unit (word or subword). Comp. indicates whether the edit is compressed. Subset specifies whether the edits capture all errors, punctuation-only errors (Pnx), or non-punctuation errors (NoPnx). NoPnx models are evaluated after removing punctuation, while Pnx models are evaluated on a version of the Dev set where all non-punctuation errors are corrected. Pruning experiments were conducted using the top two models (AraBERTv02 and CAMeLBERT), while punctuation segregation experiments used the best model (AraBERTv02). Best All results are in bold; best NoPnx and Pnx results are underlined.
@@ Appendix D LLMs Results
Table 11: LLMs results on MSA and DA GEC on the Dev sets of QALB-2014, ZAEBUC, and MADAR CODA. P-Lang is the prompt language either in English (EN) or Arabic (AR). Best average F0.5 results for each LLM are underlined; best overall results are in bold.
Appendix E MSA GEC Results
QALB-2014 ZAEBUC
P R F1 F0.5$\dagger$ P R F1 F0.5$\dagger$
AraBART 83.2 64.9 72.9 78.7$\dagger$ 87.3 70.6 78.1 83.4$\dagger$
AraT5+Morph+GED43 83.1 67.9 74.7 79.6$\dagger$ 85.2 71.2 77.6 82.0$\dagger$
AraBART+Morph+GED13 83.9 65.7 73.7 79.5$\dagger$ 87.6 73.9 80.2 84.5$\dagger$
GPT-3.5-turbo 68.6 58.6 63.2 66.3$\dagger$ 71.0 63.5 67.1 69.4$\dagger$
GPT-4o 80.7 65.7 72.4 77.2$\dagger$ 86.5 76.8 81.3 84.3$\dagger$
Fanar 69.7 63.7 66.6 68.4$\dagger$ 76.3 73.6 74.9 75.8$\dagger$
Jais-13B-Chat 49.1 36.9 42.1 46.0$\dagger$ 50.2 19.7 28.3 38.4$\dagger$
sweet 81.8 68.8 74.7 78.8$\dagger$ 85.8 72.3 78.4 82.7$\dagger$
$\textsc{sweet}^{2}$ 81.9 70.4 75.7 79.3$\dagger$ 85.8 73.3 79.1 83.0$\dagger$
$\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ 83.7 68.8 75.6 80.3$\dagger$ 86.7 73.9 79.8 83.8$\dagger$
3-Ensemble 84.9 68.8 76.0 81.1$\dagger$ 89.6 72.8 80.3 85.6$\dagger$
4-Ensemble 89.1 61.6 72.8 81.8$\ddagger$ 93.3 68.3 78.9 86.9$\dagger$
Table 12: MSA GEC results on the Dev sets of QALB-2014 and ZAEBUC. Best non-ensemble results are underlined; best overall results are in bold. $\dagger$ denotes statistical significance over the best baseline; $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
QALB-2014 QALB-2015 ZAEBUC
P R F1 F0.5$\dagger$ P R F1 F0.5$\dagger$ P R F1 F0.5$\dagger$
AraBART 84.0 64.7 73.1 79.3$\dagger$ 82.0 71.7 76.5 79.7$\dagger$ 86.0 71.6 78.2 82.7$\dagger$
AraBART+GED43 84.2 65.4 73.6 79.6$\dagger$ 81.2 72.4 76.5 79.3$\dagger$ 85.4 72.6 78.5 82.5$\dagger$
AraBART+Morph+GED43 83.9 65.7 73.7 79.5$\dagger$ 82.6 72.1 77.0 80.3$\dagger$ 85.4 73.7 79.1 82.7$\dagger$
AraBART+GED13 84.1 65.0 73.3 79.4$\dagger$ 81.5 72.7 76.8 79.5$\dagger$ 85.9 73.4 79.2 83.1$\dagger$
GPT-4o 81.5 65.5 72.6 77.7$\dagger$ 81.1 74.3 77.5 79.6$\dagger$ 84.4 75.9 79.9 82.5$\dagger$
$\textsc{sweet}^{2}$ 82.6 69.5 75.5 79.6$\dagger$ 80.0 74.3 77.0 78.8$\dagger$ 85.5 74.4 79.6 83.0$\dagger$
$\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ 84.5 67.7 75.2 80.5$\dagger$ 82.2 73.6 77.7 80.3$\dagger$ 85.7 74.1 79.5 83.1$\dagger$
3-Ensemble 85.7 67.4 75.4 81.3$\dagger$ 83.7 73.3 78.1 81.3$\dagger$ 89.7 73.7 80.9 85.9$\dagger$
4-Ensemble 89.7 60.2 72.0 81.7$\ddagger$ 88.3 66.7 76.0 82.9$\ddagger$ 93.4 68.9 79.3 87.2$\ddagger$
Table 13: MSA GEC results on the Test sets of QALB-2014, QALB-2015 (L1), and ZAEBUC. Best non-ensemble results are underlined; best overall results are in bold. $\dagger$ denotes statistical significance over the best baseline; $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
Appendix F DA GEC Results
P R F1 F0.5$\dagger$
AraT5 86.8 77.4 81.8 84.7$\dagger$
AraT5+City 87.6 79.3 83.3 85.8$\dagger$
GPT-3.5-turbo 35.5 29.7 32.3 34.1$\dagger$
GPT-4o 53.7 54.4 54.1 53.8$\dagger$
Fanar 24.5 28.8 26.4 25.2$\dagger$
Jais-13B-Chat 14.1 15.0 14.5 14.3$\dagger$
sweet 89.1 75.5 81.7 86.0$\dagger$
$\textsc{sweet}^{\text{2}}$ 87.5 73.5 79.9 84.3$\dagger$
3-Ensemble 91.7 77.4 83.9 88.4$\dagger$
4-Ensemble 93.8 72.5 81.8 88.6$\ddagger$
Table 14: DA GEC results on the MADAR CODA Dev set. Best non-ensemble results are underlined; best overall results are in bold. $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
P R F1 F0.5$\dagger$
AraT5 87.3 78.0 82.4 85.2$\dagger$
AraT5+DA Phrase 88.4 79.0 83.4 86.3$\dagger$
GPT-4o 56.1 54.8 55.5 55.9$\dagger$
sweet 89.4 76.6 82.5 86.5$\dagger$
3-Ensemble 92.2 77.7 84.3 88.9$\dagger$
4-Ensemble 94.0 72.9 82.1 88.8$\ddagger$
Table 15: DA GEC results on the Test set of MADAR CODA. Best non-ensemble results are underlined; best overall results are in bold. $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
Appendix G Error Type Statistics
QALB-2014 ZAEBUC MADAR CODA
Train Dev Test Train Dev Test Train Dev Test
Delete 6,442 346 540 305 64 66 35 0 1
Merge-B 15,063 797 795 849 180 133 404 102 95
Merge-I 15,296 812 807 851 180 133 429 109 103
M 1,466 69 63 137 32 28 159 42 56
M+O 243 17 15 7 1 8 0 0 0
O 141,752 7,380 6,976 3,203 695 792 5,604 1,179 1,193
O+X 323 24 18 20 0 4 0 0 0
P 11,379 598 687 237 51 36 18 10 9
S 5,436 247 252 169 36 51 408 77 64
X 13,592 809 743 528 110 113 911 166 151
Split 7,828 432 399 49 10 10 279 58 69
UNK 6,835 331 300 361 78 61 2,235 430 649
C 795,510 41,875 39,690 18,411 3,839 3,683 29,285 6,627 6,456
1,021,165 53,737 51,285 25,127 5,276 5,118 39,767 8,800 8,846
Table 16: Distribution of error types in QALB-2014, ZAEBUC, and MADAR CODA. UNK refers to unknown error types; C refers to correct words.
Appendix H Prompts